8x-6y=-96
2x+3y=12

Answer by stanbon(75887)

8x-6y=-96
2x+3y=12
============
8x-6y = -96
4x+6y = 24
----------------------
12x = -72
x = -6
------
Solve for "y"::
3y = -2x+12
3y = 24
y = 8
-------------
Cheers,
Stan H.
----------